11.(20) A solid cylinder of radius R, starts from rest atop a 4m high incline. The cylinder rolls without sliding. Using energy principles, determine the velocity at the bottom of the incline. 4m
Answer #1

moment of inrtia of solid cylinder = I
I = (m*R^2)/2

use,
intial potential energy = rotational kinetic energy
m*g*h = 0.5*I*w^2
m*9.8*4 = 0.5*{(m*R^2)/2}*w^2
39.2 = 0.25*R^2*w^2
w^2 = 156.8/R^2
w = 12.5/R

velocity = R*w
= R*(12.5/R)
= 12.5 m/s

Answer: 12.5 m/s